OpenStudy (anonymous):

Solve (i+1)^21 without a calculator. I know the answer, but forgot how to do it.

OpenStudy (michele_laino):

hint: we can write this: \[\Large 1 + i = \sqrt 2 {e^{i\pi /4}}\]

OpenStudy (anonymous):

Forgot to note: This question is supposed to be for Algebra II level math, I don't recall using that when I did it

OpenStudy (michele_laino):

sorry!

OpenStudy (anonymous):

My mistake :)

OpenStudy (michele_laino):

sincerely I don't know how to get the requested value without using the polar form of a complex number

OpenStudy (anonymous):

That's alright, thank you anyway

OpenStudy (michele_laino):

ok! :)

OpenStudy (triciaal):